Attendees: Bruce Glagola, Wendy VanWingeren, Dean Wyncott, and Patti Pedergnana
- Sectors 5 and 10 will be inspected with Nena at 1400 tomorrow.
- The laboratory was criticized by DOE by the central ESQ people not being out in the field with the different divisions. The APS ES&H coordinators will be reorganized into ESQ.
- Zone F’s walk-through was completed today. Zones A-E will be done tomorrow.
- There is one radioactive sample experiment ESAF submitted for the start of the run at HP-CAT.
- Dean has started sending Routine Radiation Survey emails.
- Dean is working on DEEI inspections at 35-ID.
- Dean will touch base with Central Shops on the building of the 33-ID-A deck.
- Tim Clute adjusted the air cushion on the 16-BM-D door and it closes smoothly now.
- Nancy Lazarz is paying for repair of the 13-ID-A oxygen deficiency monitor instead of ordering a new model.
- Kevin Stoll will be working on tie-in of the Sector 1 fire alarm panel into the ANL-FD panel. He has drawings and system information which was provided by A. Mashayekhi or found in the alarm panel.